The sum of the ages of Ira and Sofia is 26. The difference between 4 times Sofia's age and twice Ira's age is 8.
Answer by longjonsilver(2297)

let ages be x and y

we then have
x+y = 26 --eqn1
and
4x-2y = 8 --eqn2

eqn1, scaled up, is: 4x+4y=104

So, we have the following:
4x+4y=104
4x-2y = 8

Now subtract them, to give 6y = 96

--> y = 96/6
so y =16

from eqn1, x+y = 26, so x+16 = 26 --> x = 10

check in eqn2:
4x-2y
4*10 - 2*16
40 - 32
8.. correct

jon.